OpenStudy (anonymous):

the length of a rectangle is 4 in longer than its width. If the perimeter of the rectangle is 32 in. find its area

hartnn (hartnn):

hi again :) so, L=W+4 P =2(L+W) = 32 find L and W from here ?

OpenStudy (anonymous):

w=6 l=10

OpenStudy (anonymous):

area is 60?

hartnn (hartnn):

good! you are good at this now :)